OpenStudy (anonymous):

A nickle, a dime,and a penny are tossed. what is the probability that they do not all land the same way-they are not all heads or tails?

OpenStudy (anonymous):

Calculate the probability of all heads and the probability of all tails. Then, subtract those probabilities from 1. P(all heads) = 1/2 * 1/2 * 1/2 = 1/8 P(all tails) = 1/2 *1/2 * 1/2 = 1/8 P(not all heads or tails)= 1 - 1/8 - 1/8 = 6 / 8 = 3/4